Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (anonymous):

Please help!! i will fan and medal! A. Describe the connection of the pattern of outcomes to Pascal’s triangle. B. Look at the increase in values. What is the pattern in the number of total possible outcomes?

OpenStudy (anonymous):

If you toss a coin once, the possible results are heads (H) or tails (T). The probability of getting a H is 1/2 = 50% If you toss a coin two times, the possible results are HH, HT, TH, and TT. The probability of getting one heads and one tails is 2 out of 4 results or 2/4 = 50%

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!